OpenStudy (anonymous):

emily bought a magazine for $4 and 5 erasers.she spent a total of $29.how much did each eraser cost?

OpenStudy (anonymous):

Dang, those are some expensive erasers!

OpenStudy (anonymous):

hahaha yea

OpenStudy (anonymous):

Subtract the cost of the magazine from the 29$ an you'll see.

OpenStudy (anonymous):

5 bucks each eraser is 5 dollars

OpenStudy (anonymous):

because you take 29-4 = 25 /5 = 5 dollars